OpenStudy (solomonzelman):

how do you find the y and the x in the following equations: y+3x=5 x-y=1

OpenStudy (solomonzelman):

Well in x-y=1 you can bring y to the other side, which forms.... x=1+y lets substitute y+1 for the x in the first equation. y+3(y+1)=5 y+3y+1=5 4y+1=5 bring 1 to the other side... 4y=4 so y=1